OpenStudy (anonymous):

You are reading a book 560 pages long.You have read about 11/20 or the book.What percent of the book have you read?

OpenStudy (jamesj):

If you've x/100 of the book, you've read x% So for what x is \[ \frac{11}{20} = \frac{x}{100} \]

OpenStudy (jamesj):

Hint: \[ \frac{11}{20} = \frac{ 5 \times 11}{5 \times 20} \]

OpenStudy (jamesj):

For example, suppose you've read 1/2 of the book, what percentage is that? Well, it's x% where \( 1/2 = x/100 \). \[ \frac{1}{2} = \frac{50 \times 1}{50 \times 2} = \frac{50}{100} \] Hence 1/2 is 50% of the book.

OpenStudy (anonymous):

thank you james